# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v1

@@ 4.1.1. Experimental Data
From each of the screenshot datasets described in Section 3.1.1 (GPSCap, Screen2Words, and Clarity), we randomly selected 1000 screenshots for validation, 1000 screenshots for test, and the rest for model training, as shown in Table 3. Similar to the approach suggested by Wang et al. (Wang et al., 2021), we split the data to not share the screenshots from the same app across different splits. That is, all the apps and screenshots in the test and validation set were completely unseen during the training. This arrangement allows us to assess how well our model generalizes to previously unseen screenshots from unseen apps during the test phase.
Table 3. Evaluation dataset statistics
@@
4.1.2. UIClip Training Details
We trained the ”openai/clip-vit-base-patch32” (OpenAI, 2021b) checkpoint on the training set for a total of 5 epochs. During the training, a mini-batch gradient descent approach was employed with a batch size of 128. To optimize the learning process, the AdamW optimizer was utilized with an initial learning rate set to $5e^{-5}$.
@@ 4.1.3. Baselines
As highlighted by Bernal-Cardenas et al. (Bernal-Cardenas et al., 2019) and by Kolthoff et al. (Kolthoff et al., 2023), existing approaches perform their queries based on GUI metadata, which is absent in screenshots collected by processing app introduction images from Google Play. Consequently, we implemented a text-only retrieval approach by ourselves. In addition, we compared our model with (a) the original CLIP model without further tuning and (b) with the CLIP model fine-tuned without our GPSCap data. This resuts in three baselines:
  • OCR + BGE (Text Only). As discussed in the Section 2.4.1, PaddleOCR (Li et al., 2022b) is an industrial grade OCR library. BGE (Xiao et al., 2023) is a state-of-the-art text embedding model, trained in three steps: including pre-training with plain texts, contrastive learning on text pair dataset, and task-specific fine-tuning. First, the screenshots are embedded by using OCR + BGE in three steps: 1) the texts displayed on the GUI image are extracted with PaddleOCR; 2) the extracted text is then concatenated into a sentence, separated by semicolons; 3) the sentence is then embedded by BGE. Then, the captions are directly embedded using the BGE model.
  • CLIP. The screenshots are embedded by the image encoder of CLIP and the captions are embedded by the text encoder of CLIP. The ”openai/clip-vit-base-patch32” checkpoint is used for this evaluation.
  • UIClip-CS. The UIClip-CS is a fine-tuned model from CLIP model using the Clarity and Screen2Words dataset. The training steps is exactly same as the training of UIClip, except that our GPSCap dataset is excluded during training process. The image encoder of UIClip-CS embeds the screenshots, while its text encoder embeds the captions.
@@ 4.1.4. Evaluation Metric
4.2. Expt2: Manual Evaluation of Search Engines
A GUI search engine accepts textual queries as input and identifies the corresponding images within the GUI repository as output. Expt1 focuses on evaluating the performance of retrieval models on the test dataset. Thereby, the main limitation is the use of the screenshot-caption pairs as ground truth, that is the assumption of a direct one-to-one correspondence between a single caption (search query) and a relevant screenshot. However, such assumption contradicts real-world scenarios where one query can pertain to multiple screenshots.
To address this limitation, experiment 2 aims to evaluate different search engines, which are composed of various retrieval models and GUI repositories, through a manual evaluation. Specifically, we measure the proportion and rank of images retrieved by the GUI search engines that are deemed relevant to the query entered by the user. Through comparative analysis of the usefulness between GUing and baseline approaches, we are able to address RQ2.
@@ 4.2.1. Baselines
To answer RQ2, we undertook a comparative analysis of our search engine with two baselines, RaWi and GUing-CS.
  • RaWi. RaWi (Kolthoff et al., 2023) is a data-driven GUI prototyping approach that retrieves screenshots for reuse from Rico dataset based on natural language searches. The search engine of RaWi is a BERT-based Learning-To-Rank (LTR) model that is trained on GUI relevance data collected from a crowd-sourcing task. As Rico dataset provides the metadata of the screenshots, RaWi creates the textual representation of the screenshots by assimilating elements such as the activity names, UI component text, resource ids, and icon ids. This integrated textual representation is then compared with the user text query, forming the basis for GUI ranking.
  • GUing-CS. The architecture of GUing-CS replicates precisely that of GUing. However, the retrieval model of GUing-CS is UIClip-CS, which only uses the Clarity and Screen2Words datasets for model training. That is we explicitly remove our GPSCap dataset from the training process. Furthermore, the GUI repository for GUing-CS is comprised exclusively of the Rico and ReDraw datasets, with no utilization of the GPSRepo dataset.
@@ 4.2.2. Evaluation Method
To compare between the various search engines, we developed a tool as shown on Figure 6 The tool accepts textual queries as input and produces the top-10 screenshots from the three search engines: GUing, GUing-CS, and RaWi. The tool mixes and randomly shuffles the resulting 30 screenshots to prevent the possible bias of the evaluators. That is the origin of each search results is unclear and cannot be guessed by the evaluator.
During the assessment, evaluators are asked to perform search with the provided queries, select all screenshots they consider relevant to the query and confirm their selections via the submission button. A screenshot is deemed relevant to a query if it either displays a UI implementation of the feature delineated in the query or if it can serve as a reference or inspiration for the development of the feature.
@@
The evaluation queries were derived from the articles of AttractGroup and BuildFire, which are two companies specialized in mobile app development services and have substantial knowledge in this field. AttractGroup enumerated 138 mobile app features to be considered during mobile app development, it consists of 45 general features and 93 domain specific features spanning across 11 domain (such as E-learning, mHealth, FinTech, etc.) (Terekhov, 2023). BuildFire proposed 50 innovative app ideas for 2024, forming a good source of inspiration for prospective entrepreneurs (Hughart, 2023). From these two resources, a total of 188 mobile app features were collected. These were then divided into two sets: one set consisting of the 45 general and 50 innovative features, and another set comprising 93 domain-specific features.
A total of four evaluators took part in this assessment, each holding a master’s degree in computer science and possessing a minimum of five years’ experience in software development. Two of them performed the evaluation with the first set of queries, the other two used the second set of queries, which ensures that each query has been evaluated by two evaluators. The evaluators performed the searching, clicking and submitting for their respective queries, as described above.
@@ 4.2.3. Evaluation Metric
To measure the performance of the GUI search engine, we applied well-known information retrieval metrics (similar to Kolthoff et al. (Kolthoff et al., 2023)), including Precision@k, Mean Reciprocal Rank, and HITS@k.
@@
  • Mean Reciprocal Rank (MRR) is a statistic measure for evaluating any process that produces an ordered list of possible responses to a sample of queries. $rank_{i}$ signifies the rank of the highest ranked relevant screenshot of the i-th query and $Q$ stands for the total number of queries.
@@ 5. Evaluation Results
UIClip-CS 0.073 0.146 0.198 0.278 0.515 0.648
UIClip 0.075 0.150 0.207 0.287 0.533 0.667
5.1. RQ1 Performance of UIClip
Table 4 shows the evaluation results for the text-to-GUI retrieval task using UIClip (last row) and the baseline models on three distinct datasets.
@@ 5.1. RQ1 Performance of UIClip
Compared to the retrieval by chance, which has a recall@10 of 0.01, the CLIP model proves to be more effective in text-to-GUI retrieval, with a recall@10 exceeding 0.2 across the three datasets. This implies that, given a caption, the CLIP model has a 20 percent probability that the corresponding screenshot lies within the top 10 retrieved screenshots. Nevertheless, the UIClip model consistently outperforms the CLIP model in almost all metrics spanning the three datasets. This indicates that our fine-tuned CLIP model, UIClIp, is more adept at bridging the semantic gap between the caption and the screenshot.
@@
UIClip-CS were trained merely on the training set of Screen2Words and Clarity. Notably, UIClip delivers better performance than UIClip-CS on the Screen2Words and Clarity datasets, registering a performance improvement ranging between 0.1 to 0.3 across various metrics. This performance distinction is further accentuated in the GPSCap dataset, where UIClip-CS exhibits an inferior performance, even in comparison to the CLIP model. This anomalous outcome is hypothesized to derive from caption style variances. Specifically, the captions associated with the Screen2Words and Clarity datasets are characterized by their significant length and detail, which may possibly impede UIClip-CS’s capacity to generalize effectively on the GPSCap dataset. On a holistic level, the results of UIClip suggest that fine-tuning on our GPSCap dataset contributes substantially to the performance of CLIP models for the text-to-GUI retrieval task. This dataset enhances the performance of the CLIP model not only on the GPSCap dataset, but it also leads to performance gains on other screenshot-caption datasets.
@@
As for the text-only approach, when analysing the GPSCap and Screen2Words datasets, it becomes evident that UIClip surpasses the performance of OCR+BGE. Interestingly, when the comparison is carried out on the Clarity datasets, UIClip exhibits a deterioration in comparison to text-only approach within the given metrics recall@1, Recall@3, recall@5 and recall@10. This discrepancy can presumably be attributed to the variability in caption styles. The majority of captions in the GPSCap dataset are simple features, such as “Track your health trends” or “Custom color themes”. And the captions found within the Screen2Words dataset generally encompass high-level descriptions of the screenshot, examples being “screen showing settings options” and a “display of news articles in a news app”. In contrast, the Clarity dataset captions offer substantially more details about the screenshots, like “at the bottom left there is a skip button for users to skip the introduction”, and “in the middle of the screen a popup is displayed with a label called set date”. These detailed descriptions can easily be matched with the text displayed on the screenshots, thereby improving the performance of text-only approaches. Nevertheless, UIClip continues to surpass OCR+BGE in both recall@50 and recall@100.
@@
Overall, our findings underscore that UIClip demonstrates a superior performance than the baseline models in text-to-GUI retrieval, particularly when the goal is to search descriptions of the overall functionality and features implemented in the screens rather than verbose, documentation-like descriptions of the UI elements in the screens.
5.2. RQ2 Relevance Search Results
Table 5 shows the results of the manual evaluation for GUing and the baseline search engines. The results show that GUing consistently outperforms RaWi across all evaluation metrics. P@1 and HIT@1 show that there is 0.372 probability that GUing will return a relevant screenshot as the first result, a considerable improvement over RaWi’s rate of about 0.291. The P@10 value, indicative of the proportion of relevance results of the top-10 results returned by GUing, stands at 0.343, compared with RaWi’s rate of 0.214. The HIT@10 evaluation metric shows that in more than 91% of cases, GUing can return at least one pertinent screenshot among the top-10 results, far exceeding the percentage for RaWi, which is recorded at 70%. This is particularly interesting for UI ideation tasks, as it seems reasonable for a designer to check 10 screen suggestions for identifying at least one relevant screen (similar to screening the first result page of a google search). Finally, the MRR metric reveals that the ranking of the first relevant result returned by GUing is superior to that of RaWi.
A comparative analysis of the evaluation outcomes for GUing and GUing-CS shows a significant outperformance of GUing. This can be quantified by notable gaps in the following metrics: a gap exceeding 0.2 in regards to P@k, approximately 0.3 for HIT@k, and 0.25 in terms of MRR. The architecture of the search engines GUing and GUing-CS is completely analogous, with their only difference being the vision-language models’ training set and the repository used for searches. GUing-CS does not incorporate our GPSCap and GPSRepo datasets for its training and search operations. The large gap between GUing and GUing-CS serves as further validation of the critical importance of our GPSCap and GPSRepo datasets.
@@ 5.2. RQ2 Relevance Search Results
We carried out additional analyses on GUing and RaWi using a few queries in which the original search engine of the screenshots was not concealed. The example of the returned screenshots of GUing and RaWi with the query ”sleeping track” are shown on Figure 8 and 8. Upon examination, it became apparent that the stylistic differences between screenshots returned by two search engines were considerable. This difference lays on three aspects:
  • The shape of the screenshots returned by GUing is less neat. A large portion of screenshots within the GPSRepo dataset is derived from the surrounded screenshots, in which the screenshot area varies in shape across diverse images. Even the shape is less neat, these screenshots can still be used for the inspiration of GUI design.
  • GUing returns a wider variety of screenshots than RaWi. As shown on Figure 8, the presence of empty or settings pages among the screenshots returned by RaWi is thought to be an inherent limitation of the text-only retrieval approach, as they only offer minor contribution to sparking creativity in GUI design. Text-only retrieval approach compares the similarity of the query with text or metadata found on each screenshot. As setting pages typically contain an abundance of text, the likelihood of heightened query similarity is increased. Empty pages, on the other hand, are likely a consequence of the Rico data collection method, which involves the automatic exploration of apps. When performing automatic exploration, there is no initial data in the app, such as sleep events record.
  • The modernity of the screenshots returned by GUing vastly exceeds those resulting from RaWi. This discrepancy can feasibly be attributed to the different collection periods of the underlying datasets. The GPSRepo dataset was collected recently, whereas the Rico dataset was created in 2017. The GPSRepo dataset can be easily updated by processing introduction images of new apps when added to the Google Play store.
Figure 7. Top-10 screenshots returned by GUing with query ”sleep tracking”
@@
Top-10 screenshots returned by GUing with query ”sleep tracking”
@@
6. Exploring UIClip as Foundation Model for GUI Tasks
In the previous section, we evaluated the performance of UIClip (as a main component of GUing) for the text-to-GUI retrieval task. However, as a tuned foundation model, UIClip can be adapted for various other tasks. In this section, we discuss the broader application of UIClip to GUI-related tasks, including GUI classification, sketch-to-GUI retrieval, GUI captioning, and GUI generation—along with a preliminary evaluation.

### v2

@@ 4.1.1. Experimental Data
From each of the screenshot datasets described in Section 3.1.1 (SCapRepo, Screen2Words, and Clarity), we randomly selected 1000 screenshots for validation, 1000 screenshots for test, and the rest for model training, as shown in Table 3. Similar to the approach suggested by Wang et al. (Wang et al., 2021), we split the data to not share the screenshots from the same app across different splits. That is, all the apps and screenshots in the test and validation set were completely unseen during the training. This arrangement allows us to assess how well our model generalises to previously unseen screenshots from unseen apps during the test phase.
Table 3. Statistics of the evaluation datasets.
@@
4.1.2. GUIClip Training Details
We trained the ”openai/clip-vit-base-patch32” (OpenAI, 2021b) checkpoint on the training set for a total of 5 epochs. During the training, a mini-batch gradient descent approach was employed with a batch size of 128. To optimise the learning process, the AdamW optimiser was utilised with an initial learning rate set to $5e^{-5}$.
@@ 4.1.3. Baselines
As highlighted by Bernal-Cardenas et al. (Bernal-Cardenas et al., 2019) and by Kolthoff et al. (Kolthoff et al., 2023), existing approaches perform their queries based on GUI metadata, which is absent in screenshots collected by processing app introduction images from Google Play. Consequently, we implemented a text-only retrieval approach by ourselves. In addition, we compared our model with (a) the original CLIP model without further fine-tuning and (b) with the CLIP model fine-tuned without our SCapRepo data. This results in three baselines:
  • OCR + BGE (Text Only): As discussed in the Section 2.4.1, PaddleOCR (Li et al., 2022b) is an industrial grade OCR library. BGE (Xiao et al., 2023) is a state-of-the-art text embedding model, trained in three steps: pre-training with plain texts, contrastive learning on text pair dataset, and task-specific fine-tuning. First, the screenshots are embedded by using OCR + BGE in three steps: 1) texts displayed on the GUI images are extracted with PaddleOCR; 2) the extracted text is then concatenated into a sentence, separated by semicolons; 3) the sentence is then embedded by BGE. Finally, the captions are directly embedded using the BGE model.
  • CLIP: The screenshots are embedded by the image encoder of CLIP and the captions are embedded by the text encoder of CLIP. The ”openai/clip-vit-base-patch32” checkpoint is used for this evaluation.
  • GUIClip-CS: This is a fine-tuned model from CLIP using the Clarity and Screen2Words dataset. The training process is exactly the same as for GUIClip, except that our SCapRepo dataset is excluded during the training. The image encoder of GUIClip-CS embeds the screenshots, while its text encoder embeds the captions.
@@ 4.1.4. Evaluation Metric
4.2. Exp2: Manual Evaluation of Search Engines
A GUI search engine accepts textual queries as input and identifies the corresponding images within the GUI repository as output. Exp1 focuses on evaluating the performance of retrieval models on the test datasets. Thereby, the main limitation is the use of the screenshot-caption pairs as ground truth, assuming a direct one-to-one correspondence between a single caption (search query) and a relevant screenshot. However, such assumption contradicts real-world scenarios where one query can pertain to multiple screenshots.
To address this limitation, our second experiment aims to evaluate different search engines composed of various retrieval models and GUI repositories, through a manual assessment. Specifically, we measure the proportion and ranks of images retrieved by the GUI search engines that are relevant to the query entered by the user. Through a comparative analysis of the results usefulness between GUing and baseline approaches, we are then able to address RQ2.
@@ 4.2.1. Baselines
To answer RQ2, we compared our search engine to two baselines: RaWi and GUing-CS.
  • RaWi. RaWi (Kolthoff et al., 2023) is a data-driven GUI prototyping approach that retrieves screenshots for reuse from Rico dataset based on natural language searches. The search engine of RaWi is a BERT-based (Devlin et al., 2019) Learning-To-Rank (LTR) model that is trained on GUI relevance data collected from crowd-workers. As Rico dataset provides the metadata of the screenshots, RaWi creates the textual representation of the screenshots by assimilating elements such as the activity names, UI component text, resource ids, and icon ids. This integrated textual representation is then compared with the user text query, forming the basis for the GUI ranking.
  • GUing-CS. The architecture of GUing-CS replicates that of GUing. However, the retrieval model of GUing-CS is GUIClip-CS, which only uses the Clarity and Screen2Words datasets for training. That is, we explicitly remove our SCapRepo dataset from the training process. Furthermore, the GUI repository for GUing-CS is comprised exclusively of the Rico and ReDraw datasets, without the ScreenRepo dataset.
@@ 4.2.2. Evaluation Method
To compare the search engines, we developed a tool as shown on Figure 6. The tool accepts textual queries as input and produces the top-10 screenshots from the three search engines: GUing, GUing-CS, and RaWi. The tool mixes and randomly shuffles the resulting 30 screenshots to prevent possible biases by the evaluators. The origin of each search results is unclear and cannot be guessed by the evaluator. As shown on Figure 6, GUIs might be displayed multiple times in the result due to their retrieval by more than one search engine.
During the assessment, the evaluators are asked to search for provided text queries, select all screenshots they consider relevant and confirm their selections via the submission button. A screenshot is relevant to a query if it either displays a UI implementation of the feature delineated in the query or if it can serve as a reference or inspiration for the development of that feature.
@@
The evaluation queries were derived from articles by AttractGroup and BuildFire, which are two companies specialised in mobile app development services and have substantial knowledge in this field. AttractGroup’s article enumerates 138 mobile app features to be considered during mobile app development, consisting of 45 general features and 93 domain specific features spanning across 11 domains (such as E-learning, mHealth, FinTech, etc.) (Terekhov, 2023). BuildFire’s article proposes 50 innovative app ideas for 2024, forming a good source of inspiration for prospective entrepreneurs (Hughart, 2023). From these two resources, we collected a total of 188 mobile app features. These were then divided into two sets: one set consisting of the 45 general and 50 innovative features, and another set comprising 93 domain-specific features.
Four evaluators took part in this experiment, each holding a master degree in computer science and having a minimum of five years of software development experience. Two of them performed the evaluation with the first set of queries, the other two used the second set of queries. This ensured that each query was evaluated independently by two evaluators. The evaluators performed the search and submitted for their respective results as described above.
@@ 4.2.3. Evaluation Metric
To measure the performance of the GUI search engine, we applied the common information retrieval metrics Precision@k, Mean Reciprocal Rank, and HITS@k (similar to Kolthoff et al. (Kolthoff et al., 2023)):
@@
  • Mean Reciprocal Rank (MRR) is a measure for evaluating any process that produces an ordered list of possible responses to a sample of queries. $rank_{i}$ signifies the rank of the highest ranked relevant screenshot of the i-th query. $Q$ stands for the total number of queries.
@@ 5. Evaluation Results
GUIClip-CS 0.073 0.146 0.198 0.278 0.515 0.648
GUIClip 0.075 0.150 0.207 0.287 0.533 0.667
5.1. RQ1 Performance of GUIClip
@@
Table 4 shows the evaluation results for the text-to-GUI retrieval task using GUIClip (last row) and the baseline models on three distinct datasets. Compared to the retrieval by chance, which has a recall@10 of 0.01, the CLIP model proves to be more effective in text-to-GUI retrieval, with a recall@10 exceeding 0.2 across the three datasets. This implies that, given a caption, the CLIP model has a 20 percent probability that the corresponding screenshot lies within the top 10 retrieved screenshots. Nevertheless, the GUIClip model consistently outperforms the CLIP model in almost all metrics and for the three datasets. This indicates that our fine-tuned CLIP model, GUIClip, is more adept at bridging the semantic gap between the caption and the screenshot.
@@
GUIClip-CS were trained merely on the Screen2Words and Clarity datasets. Notably, GUIClip delivers better performance than GUIClip-CS on the Screen2Words and Clarity datasets, registering a performance improvement ranging between 0.1 to 0.3 across various metrics. This performance distinction is further accentuated in the SCapRepo dataset, where GUIClip-CS exhibits an inferior performance, even when compared to the CLIP model. This outcome is hypothesised to derive from caption style variances. Specifically, the captions associated with the Screen2Words and Clarity datasets are significantly lengthy and detailed, which may possibly impede GUIClip-CS’s capacity to generalise effectively on the SCapRepo dataset. On a holistic level, the results of GUIClip suggest that fine-tuning on our SCapRepo dataset contributes substantially to the performance of CLIP models for the text-to-GUI retrieval task. This dataset enhances the performance of the CLIP model not only on the SCapRepo dataset, but also leads to performance gains on other screenshot-caption datasets.
@@
As for the text-only approach, when analysing the SCapRepo and Screen2Words datasets, it becomes evident that GUIClip surpasses the performance of OCR+BGE. Interestingly, when the comparison is carried out on the Clarity datasets, GUIClip exhibits a deterioration in comparison to text-only approach for recall@1, Recall@3, recall@5 and recall@10. This discrepancy can presumably be attributed to the variability in caption styles. The majority of captions in the SCapRepo dataset are app features such as “Track your health trends” or “Custom colour themes”. And the captions found within the Screen2Words dataset generally encompass high-level descriptions of the screenshot, such as “screen showing settings options” and a “display of news articles in a news app”. In contrast, the Clarity dataset captions offer substantially more details about the screenshots, like “at the bottom left there is a skip button for users to skip the introduction”, and “in the middle of the screen a popup is displayed with a label called set date”. These detailed descriptions can easily be matched with the text displayed on the screenshots, thereby improving the performance of text-only approaches. Nevertheless, GUIClip continues to surpass OCR+BGE in both recall@50 and recall@100.
@@
Overall, our findings underscore that GUIClip outperforms the baseline models in text-to-GUI retrieval, particularly when the goal is to search for descriptions of an overall functionality and features implemented in the screens rather than verbose, documentation-like descriptions of the screen’s UI elements.
5.2. RQ2 Relevance of Search Results
@@
Table 5 shows the results of the manual evaluation for GUing and the baseline search engines. The results show that GUing consistently outperforms RaWi across all evaluation metrics. P@1 and HIT@1 show that there is 0.372 probability that GUing will return a relevant screenshot as the first result, a considerable improvement over RaWi’s rate of about 0.291. The P@10 value, indicative of the proportion of relevance results of the top-10 results returned by GUing, stands at 0.343, compared with RaWi’s rate of 0.214. The HIT@10 evaluation metric shows that, in more than 91% of cases, GUing can return at least one pertinent screenshot among the top-10 results, far exceeding the percentage for RaWi, which is recorded at 70%. This is particularly interesting for GUI ideation tasks, as it seems reasonable for a designer to check 10 screen suggestions for identifying at least one relevant screen (similar to screening the first result page of a google search). Finally, the MRR metric reveals that the ranking of the first relevant result returned by GUing is superior to that of RaWi.
The results also show that GUing significantly outperforms GUing-CS too. This can be quantified by notable gaps in the following metrics: a gap exceeding 0.2 for P@k, approximately 0.3 for HIT@k, and 0.25 in terms of MRR. The architectures of the search engines GUing and GUing-CS are identical, except the training set for the vision-language models and the repository used for searches. GUing-CS does not incorporate our SCapRepo and ScreenRepo datasets for its model training and search operation. The large gap between GUing and GUing-CS demonstrates the importance of our SCapRepo and ScreenRepo datasets.
Figure 7. Top-10 screenshots returned by GUing for the query ”sleep tracking”.
@@
Top-10 screenshots returned by GUing for the query ”sleep tracking”.
@@
To understand the difference between the screenshots obtained using GUing and Rawi, we carried out additional analyses on each search engine with a few queries. For example, the screenshots returned by GUing and RaWi for the query ”sleeping track” are shown on Figure 8 and 8. Upon examination, it became apparent that there are considerable stylistic differences between the screenshots returned by the two search engines, mainly due to three factors:
  • The shape of the screenshots returned by GUing is less neat. A large portion of screenshots within the ScreenRepo dataset are derived from the surrounded screenshots, in which the screenshot area varies in shape across diverse images. However, these screenshots are still useful to inspire GUI design.
  • GUing returns a wider variety of screenshots than RaWi. As shown on Figure 8, the presence of empty or settings pages among the screenshots returned by RaWi is thought to be an inherent limitation of the text-only retrieval approach, as they barely offer any contribution to sparking creativity in GUI design. Text-only retrieval approaches compare the similarity of the query with text or metadata found on each screenshot. As setting pages typically contain much text, they will likely be matched to many queries. Empty pages are likely a result of the Rico data collection method, which involves the GUI exploration of apps. When performing crowdsourced or automated exploration, the app often lack initial data, such as recorded sleep events.
  • The modernity of the screenshots returned by GUing vastly exceeds those resulting from RaWi. This discrepancy can be attributed to the collection periods of the underlying datasets. The ScreenRepo dataset was collected recently, whereas the Rico dataset was created in 2017. ScreenRepo can easily be updated by processing introduction images of new apps when added to the Google Play store.
6. Exploring GUIClip as a Foundation Model for GUI Tasks
In the previous section, we evaluated the performance of GUIClip (as a main component of GUing) for the text-to-GUI retrieval task. In this section, we address the following research question:
RQ3: Does GUIClip also show a high performance for other GUI-related tasks? The CLIP model, as a foundation model, has been extensively applied across various vision-language tasks, consistently delivering promising results (Zhang et al., 2024). As GUIClip is a fine-tuned version specifically with GUI images, it has the potential to perform well for other GUI-related tasks. Particularly, the question is whether GUIClip can also outperform CLIP in other GUI-related tasks.
To explore this, we conducted two experiments: one focusing on GUI classification and one on sketch-to-GUI retrieval.